OpenStudy (anonymous):

Using elimination, substitution solve A A. 3x - 9y = 3 6x - 3y = (-24)

OpenStudy (anonymous):

@hereweg0

OpenStudy (anonymous):

it belongs in the math section

OpenStudy (anonymous):

but okay. 3x - 9y = 3 lets get x alone. 3x=3+9y

OpenStudy (anonymous):

yeah i dont care

OpenStudy (kasim17):

1)Elimination by coefficient method 3x - 9y = 3 .....( i ) 6x - 3y = - 24 ...( ii ) Multiplying ( i ) by 2,we get 6x - 18y = 6 .... ( iii ) Subtracting ( ii ) from ( iii ),we get 6x - 18y = 6 6x - 3y = - 24 (-) (+) (+) --------------------- - 15y = 30 y = - 2 Substituting y = - 2 in ( ii ),we get 6x - 3(- 2) = - 24 6x + 6 = - 24 6x = - 30 x = - 5 x = - 5 and y = - 2 is the solution of given simultaneous equations. 2)Substitution method 3x - 9y = 3 Dividing throughout by 3,we get x - 3y = 1 x = 1 + 3y ..... ( i ) 6x - 3y = - 24 .....( ii ) Substituting ( i ) in ( ii ),we get 6(1 + 3y) - 3y = - 24 6 + 18y - 3y = - 24 15y = - 30 y = - 2 Substituting y = - 2 in ( i ),we get x = 1 + 3y x = 1 + 3(- 2) x = 1 - 6 x = - 5 x = - 5 and y = - 2 is the solution of given simultaneous equations. According to me both the methods are efficient and easy. It varies from individual to individual to decide which method is efficient. So my opinion to you is that learn both the methods and practice hard and try to master them. !!!HOPE I HELPED!!!!

OpenStudy (anonymous):

belong in math

OpenStudy (anonymous):

hmm nice haha

OpenStudy (anonymous):

this is science....

OpenStudy (anonymous):

Thanks kasim17